OpenStudy (anonymous):

Hello need help reducing this ((x^3+x/ x^Z+X)(x^z-1/X^z+1)) please show me the steps.

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

what is x^z?

OpenStudy (anonymous):

it was in the problem

OpenStudy (anonymous):

its a variable

OpenStudy (anonymous):

then this problem cannot be factored.

OpenStudy (anonymous):

the answer is x-1 but idk how that is

OpenStudy (anonymous):

if z =2, then the answer is x-1

OpenStudy (anonymous):

how did u get that?

OpenStudy (anonymous):

substitute z = 2 and solve. you will get x-1

OpenStudy (anonymous):

\[(x ^{3}+x) \div (x ^{2}+x) = x(x ^{2}+1) \div x(x+1) = (x ^{2}+1) \div (x+1)\]

OpenStudy (anonymous):

\[(x ^{2}-1) \div (x ^{2}+1) = (x-1)(x+1) \div (x ^{2}+1)\]

OpenStudy (anonymous):

multiply both, you cancel out x^2+1 and x+1, you are left with x-1

OpenStudy (anonymous):

k let me attach the file bc i think the 2 is a z

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you will get the answer x-1 only if z = 2

OpenStudy (anonymous):

does the attachment work

OpenStudy (anonymous):

looks like a 2 to me

OpenStudy (anonymous):

ok lol im going to re evaluate ur steps